Logiweb diagnose of ijcar

Up Help

Error in proof of [ [ Prop 3.2f_{2} [ [ Lemma expected [ S2 ] ] ] ] ]

The pyk compiler, version 0.grue.20060417 by Klaus Grue GRD-2006-05-26.UTC:13:59:58.555187 = MJD-53881.TAI:14:00:31.555187 = LGT-4655368831555187e-6